# RULE: Reliable Multimodal RAG for Factuality in Medical Vision Language Models

### v1

@@ Abstract
The recent emergence of Medical Large Vision Language Models (Med-LVLMs) has enhanced medical diagnosis. However, current Med-LVLMs frequently encounter factual issues, often generating responses that do not align with established medical facts. Retrieval-Augmented Generation (RAG), which utilizes external knowledge, can improve the factual accuracy of these models but introduces two major challenges. First, limited retrieved contexts might not cover all necessary information, while excessive retrieval can introduce irrelevant and inaccurate references, interfering with the model’s generation. Second, in cases where the model originally responds correctly, applying RAG can lead to an over-reliance on retrieved contexts, resulting in incorrect answers. To address these issues, we propose RULE, which consists of two components. First, we introduce a provably effective strategy for controlling factuality risk through the calibrated selection of the number of retrieved contexts. Second, based on samples where over-reliance on retrieved contexts led to errors, we curate a preference dataset to fine-tune the model, balancing its dependence on inherent knowledge and retrieved contexts for generation. We demonstrate the effectiveness of RULE on three medical VQA datasets, achieving an average improvement of 20.8% in factual accuracy. We publicly release our benchmark and code in https://github.com/richard-peng-xia/RULE.
@@ 1 Introduction
Artificial Intelligence (AI) has showcased its potential in medical diagnosis, including disease identification, treatment planning, and recommendations Tăuţan et al. (2021); Wang et al. (2019); Ye et al. (2021); Xia et al. (2024b); Li et al. (2024). In particular, the recent development of Medical Large Vision Language Models (Med-LVLMs) has introduced more accurate and customized solutions to clinical applications Li et al. (2023); Moor et al. (2023); Zhang et al. (2023); Wu et al. (2023). While Med-LVLMs have demonstrated promising performance, they remain prone to generating responses that deviate from factual information, potentially resulting in inaccurate medical diagnoses. This susceptibility to hallucination underscores the need for enhanced mechanisms to ensure factual alignment in critical medical applications (see an example in Figure 1(a)) Royer et al. (2024); Xia et al. (2024a)). Such errors pose a significant risk to clinical decision-making processes and can lead to adverse outcomes.
Recently, Retrieval-Augmented Generation (RAG) Gao et al. (2023) has emerged as a promising method for enhancing the factual accuracy of responses from Med-LVLMs. By integrating external, reliable data sources, RAG guides the model in producing factual medical responses, enriching its knowledge base with supplementary information. For example, RAG has been used in tasks such as visual question answering (VQA) Yuan et al. (2023) and report generation Kumar and Marttinen (2024); Tao et al. (2024). However, as illustrated in Figure 1(b) and Figure 1(c), directly applying RAG strategy to Med-LVLMs presents two significant challenges: (1) A small number of retrieved contexts may not cover the reference knowledge required for the question, thus limiting the model’s factual accuracy. Conversely, a large number of retrieved contexts may include low-relevance and inaccurate references, which can interfere with the model’s generation; (2) Med-LVLMs may overly rely on the retrieved information. In this situation, the model might correctly answer on its own, but incorporating the retrieved contexts could lead to incorrect responses.
@@
Our primary contributions of this paper is RULE, which introduces an innovative approach to enhance retrieval-based Med-LVLMs. RULE not only controls factual risk by calibrating the selection of reference contexts but also balances the model’s knowledge and retrieved contexts through preference fine-tuning using a curated preference dataset. Across three medical Visual Question Answering (VQA) benchmarks, including radiology and ophthalmology, our empirical results demonstrate that RULE effectively improves the factual accuracy of Med-LVLMs, achieving a 8.06% improvement over the best prior methods for mitigating hallucination. In addition, empirically verify the effectiveness of the proposed components and demonstrate the compatibility of RULE.
@@ 18       Get the predictions of the model w/ KBPT and retrieval $a\leftarrow\mathcal{M}(x,(q,T_{k}))$
Table 2: Factuality performance (%) of Med-LVLMs on the three datasets. Notably, we report the accuracy, precision, recall, and F1 score. The best results and second best results are bold and underlined, respectively.
Models IU-Xray Havard-FairVLMed MIMIC-CXR
@@ 4.1 Experimental Setups
Evaluation Datasets. To ensure that the retrieved report content is relevant to the visual question-answering content and to facilitate experimentation, we utilize three medical vision-language datasets, i.e., MIMIC-CXR Johnson et al. (2019), IU-Xray Demner-Fushman et al. (2016), and Harvard-FairVLMed Luo et al. (2024), encompassing radiology and ophthalmology. The training set is split into two parts: one part is used to train the retriever (Section 3.1), and the other part is used to construct the preference dataset for KBPT (Section 3.3).
@@
Evaluation Metrics. We use Accuracy as the primary metric and, for detailed comparisons, we also adopt Precision, Recall, and F1 Score.
@@ 4.2 Results
Comparison with Baseline Methods. We present the results of a comparison between RULE and various hallucination reduction methods in Table 2. According to these results, RULE demonstrates the best overall performance, effectively and accurately diagnosing diseases with an average accuracy improvement of 20.8% across all datasets. We also observe that RULE performs notably better on the IU-Xray and Harvard-FairVLMed compared to MIMIC-CXR. This difference is attributed to the excessive length of the reports available for retrieval in MIMIC-CXR, where overly long references tend to confuse the Med-LVLM. In addition, even when dealing with the relatively niche ophthalmology data (i.e., Harvard-FairVLMed), RULE demonstrates superior results, significantly enhancing the factual accuracy of the Med-LVLM. In contrast, the performance of decoding methods is quite unstable, showing significant rates of missed or incorrect diagnoses across different datasets, as indicated by the precision and recall values.
@@ 5 Related Work
Factuality in Med-LVLMs. The rapid development of Large Vision and Language Models (LVLMs) Liu et al. (2023b, a); Zhu et al. (2023); Alayrac et al. (2022); Zhou et al. (2024a, b) has begun to impact medical diagnosis. A series of Med-LVLMs Li et al. (2023); Moor et al. (2023); Wu et al. (2023); Zhang et al. (2023), represented by LLaVA-Med, have emerged, demonstrating impressive performance across various medical image modalities. However, Med-LVLMs still exhibit significant factual errors, producing medical responses that conflict with the visual medical information. This could potentially lead to misdiagnoses or missed diagnoses. Recently, several benchmarks Royer et al. (2024); Xia et al. (2024a) have been established to evaluate the accuracy of Med-LVLMs in tasks such as VQA or report generation. Beyond evaluating factuality, improving the factual accuracy of Med-LVLMs remains an underexplored area.
Retrieval Augmented Generation. RAG has recently been recognized as a promising solution Gao et al. (2023). It enhances the model’s ability to generate accurate facts by incorporating contextual information from external datasets. In medical multimodal analysis, the RAG approach has been applied to various tasks such as medical VQA Yuan et al. (2023) and report generation Kumar and Marttinen (2024); Tao et al. (2024); He et al. (2024). However, in Med-LVLMs, applying RAG-based approaches overlook two critical issues: the number of retrieved contexts and whether the model overly relies on these reference. These factors can significantly affect the model’s performance and may even degrade it. In RULE, we systematically address these challenges and enhance the factuality of Med-LVLMs.
@@ Appendix D Proofs
$=$ $1-\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{\mathcal{M}(x_{i},(q_{i},\{T_{k}\}_{i=1}^{N}))=y_{i}\}$
$=$ $\frac{1}{N}\sum_{i=1}^{N}(1-\mathbf{1}\{\mathcal{M}(x_{i},(q_{i},\{T_{k}\}_{i=1}^{N}))=y_{i}\})$

### v2

@@ Abstract
The recent emergence of Medical Large Vision Language Models (Med-LVLMs) has enhanced medical diagnosis. However, current Med-LVLMs frequently encounter factual issues, often generating responses that do not align with established medical facts. Retrieval-Augmented Generation (RAG), which utilizes external knowledge, can improve the factual accuracy of these models but introduces two major challenges. First, limited retrieved contexts might not cover all necessary information, while excessive retrieval can introduce irrelevant and inaccurate references, interfering with the model’s generation. Second, in cases where the model originally responds correctly, applying RAG can lead to an over-reliance on retrieved contexts, resulting in incorrect answers. To address these issues, we propose RULE, which consists of two components. First, we introduce a provably effective strategy for controlling factuality risk through the calibrated selection of the number of retrieved contexts. Second, based on samples where over-reliance on retrieved contexts led to errors, we curate a preference dataset to fine-tune the model, balancing its dependence on inherent knowledge and retrieved contexts for generation. We demonstrate the effectiveness of RULE on medical VQA and report generation tasks across three datasets, achieving an average improvement of 47.4% in factual accuracy. We publicly release our benchmark and code in https://github.com/richard-peng-xia/RULE.
@@ 1 Introduction
Artificial Intelligence (AI) has showcased its potential in medical diagnosis, including disease identification, treatment planning, and recommendations Tăuţan et al. (2021); Wang et al. (2019); Ye et al. (2021); Xia et al. (2024b); Hu et al. (2024b, a). In particular, the recent development of Medical Large Vision Language Models (Med-LVLMs) has introduced more accurate and customized solutions to clinical applications Li et al. (2023); Moor et al. (2023); Zhang et al. (2023); Wu et al. (2023). While Med-LVLMs have demonstrated promising performance, they remain prone to generating responses that deviate from factual information, potentially resulting in inaccurate medical diagnoses. This susceptibility to hallucination underscores the need for enhanced mechanisms to ensure factual alignment in critical medical applications (see an example in Figure 1(a)) Royer et al. (2024); Xia et al. (2024a)). Such errors pose a significant risk to clinical decision-making processes and can lead to adverse outcomes.
Recently, Retrieval-Augmented Generation (RAG) Gao et al. (2023); Qu et al. (2024a, b) has emerged as a promising method for enhancing the factual accuracy of responses from Med-LVLMs. By integrating external, reliable data sources, RAG guides the model in producing factual medical responses, enriching its knowledge base with supplementary information. For example, RAG has been used in tasks such as visual question answering (VQA) Yuan et al. (2023) and report generation Kumar and Marttinen (2024); Tao et al. (2024). However, as illustrated in Figure 1(b) and Figure 1(c), directly applying RAG strategy to Med-LVLMs presents two significant challenges: (1) A small number of retrieved contexts may not cover the reference knowledge required for the question, thus limiting the model’s factual accuracy. Conversely, a large number of retrieved contexts may include low-relevance and inaccurate references, which can interfere with the model’s generation; (2) Med-LVLMs may overly rely on the retrieved information. In this situation, the model might correctly answer on its own, but incorporating the retrieved contexts could lead to incorrect responses.
@@
Our primary contributions of this paper is RULE, which introduces an innovative approach to enhance retrieval-based Med-LVLMs. RULE not only controls factual risk by calibrating the selection of reference contexts but also balances the model’s knowledge and retrieved contexts through preference fine-tuning using a curated preference dataset. Across three medical Visual Question Answering (VQA) and report generation benchmarks, including radiology and ophthalmology, our empirical results demonstrate that RULE effectively improves the factual accuracy of Med-LVLMs, achieving a 14.46% improvement over the best prior methods for mitigating hallucination. In addition, empirically verify the effectiveness of the proposed components and demonstrate the compatibility of RULE.
@@ 18       Get the predictions of the model w/ KBPT and retrieval $a\leftarrow\mathcal{M}(x,(q,T_{k}))$
Table 2: Factuality performance (%) of Med-LVLMs on the three VQA datasets. Notably, we report the accuracy, precision, recall, and F1 score. The best results and second best results are bold and underlined, respectively.
Models IU-Xray Harvard-FairVLMed MIMIC-CXR
@@
Table 3: Factuality performance (%) of Med-LVLMs on the three report generation datasets. Notably, we report the average BLEU, ROUGE-L, METEOR.
Models IU-Xray MIMIC-CXR Harvard-FairVLMed
BLEU ROUGE-L METEOR BLEU ROUGE-L METEOR BLEU ROUGE-L METEOR
LLaVA-Med-1.5 9.64 12.26 8.21 12.11 13.05 11.16 18.11 11.36 10.75
+ Greedy 11.47 15.38 12.69 16.63 14.26 14.19 17.98 11.49 13.77
+ Beam Search 12.10 16.21 13.17 16.97 14.74 14.43 18.37 12.62 14.50
+ DoLa 11.79 15.82 12.72 17.11 14.89 14.81 18.26 12.51 14.51
+ OPERA 10.66 14.70 12.01 15.40 12.52 13.72 16.59 11.47 13.63
+ VCD 10.42 14.14 11.59 15.18 12.30 13.38 16.73 11.38 13.89
+ RULE (Ours) 27.53 23.16 27.99 18.61 15.96 17.42 22.35 14.93 17.74
@@ 4.1 Experimental Setups
Evaluation Datasets. To ensure that the retrieved report content is relevant to the visual question content and to facilitate experimentation, we utilize three medical vision-language datasets, i.e., MIMIC-CXR Johnson et al. (2019), IU-Xray Demner-Fushman et al. (2016), and Harvard-FairVLMed Luo et al. (2024), encompassing radiology and ophthalmology. The training set is split into two parts: one part is used to train the retriever (Section 3.1), and the other part is used to construct the preference dataset for KBPT (Section 3.3).
@@
Evaluation Metrics. For Med-VQA task, we use Accuracy as the primary metric and, for detailed comparisons, we also adopt Precision, Recall, and F1 Score. For report generation task, we use BLEU Score Papineni et al. (2002), ROUGE-L Lin (2004) and METEOR Banerjee and Lavie (2005) as the metrics.
@@ 4.2 Results
Comparison with Baseline Methods. We present the results of a comparison between RULE and various hallucination reduction methods in Table 2. According to these results, RULE demonstrates the best overall performance, effectively and accurately diagnosing diseases with an average accuracy improvement of 47.4% on two tasks across all datasets. We also observe that RULE performs notably better on the IU-Xray and Harvard-FairVLMed compared to MIMIC-CXR. This difference is attributed to the excessive length of the reports available for retrieval in MIMIC-CXR, where overly long references tend to confuse the Med-LVLM. Even when dealing with the relatively niche ophthalmology data (i.e., Harvard-FairVLMed), RULE demonstrates superior results, significantly enhancing the factual accuracy of the Med-LVLM. In contrast, the performance of decoding methods is quite unstable, showing significant rates of missed or incorrect diagnoses across different datasets, as indicated by the precision and recall values.
@@ 5 Related Work
Factuality in Med-LVLMs. The rapid development of Large Vision and Language Models (LVLMs) Liu et al. (2023b, a); Zhu et al. (2023); Alayrac et al. (2022); Zhou et al. (2024a, b); Xia et al. (2024c, 2023) has begun to impact medical diagnosis. A series of Med-LVLMs Li et al. (2023); Moor et al. (2023); Wu et al. (2023); Zhang et al. (2023), represented by LLaVA-Med, have emerged, demonstrating impressive performance across various medical image modalities. However, Med-LVLMs still exhibit significant factual errors, producing medical responses that conflict with the visual medical information Xia et al. (2024a); Su et al. (2024). This could potentially lead to misdiagnoses or missed diagnoses. Recently, several benchmarks Royer et al. (2024); Xia et al. (2024a) have been established to evaluate the accuracy of Med-LVLMs in tasks such as VQA or report generation. Beyond evaluating factuality, improving the factual accuracy of Med-LVLMs remains an underexplored area.
Retrieval Augmented Generation. RAG has recently been recognized as a promising solution Gao et al. (2023); Sun et al. (2024). It enhances the model’s ability to generate accurate facts by incorporating contextual information from external datasets. In medical multimodal analysis, the RAG approach has been applied to various tasks such as medical VQA Yuan et al. (2023) and report generation Kumar and Marttinen (2024); Tao et al. (2024); He et al. (2024). However, in Med-LVLMs, applying RAG-based approaches overlook two critical issues: the number of retrieved contexts and whether the model overly relies on these reference. These factors can significantly affect the model’s performance and may even degrade it. In RULE, we systematically address these challenges and enhance the factuality of Med-LVLMs.
@@ Appendix D Proofs
$=$ $1-\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{\mathcal{M}(x_{i},(q_{i},\{T_{k}\}_{i=1}^{N}))$
$=$ $y_{i}\}$
$=$ $\frac{1}{N}\sum_{i=1}^{N}(1-\mathbf{1}\{\mathcal{M}(x_{i},(q_{i},\{T_{k}\}_{i=1}^{N}))$
$=$ $y_{i}\})$